OpenStudy (anonymous):

Write a script for this is the real-world scenario and equation:

OpenStudy (anonymous):

Write a script that explains the use of additive and multiplicative inverses in solving a real-world problem. A bank charges a monthly fee of $5.00 for using a debit card. In addition, the bank charges $0.25 for each transaction made with the debit card. If you pay the bank a total of $15.00 one month for using your debit card, how many transactions did you make in that month? Equation: 5 + 0.25x = 15

OpenStudy (anonymous):

PLEASE DONT LEAVE I REALLLY DONT KNOW WHAT TO DO

OpenStudy (anonymous):

ok so you have created the correct equation, now you have to try and re wright that equation so that x is on one side and all the numbers are on the other side. Step 1 take away five from both sides Post what you get

OpenStudy (anonymous):

ok just a heads up I didnt wright the above ^ it was given to me i jjust had no idea what to do but ill get the fives

OpenStudy (anonymous):

when you say both sides you mean 5 + 0.25x = 15 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so think of it like this your equation is:\[5 + 0.25 = 15\] take five from both sides to get this:\[5 - 5 + 0.25x = 15 - 5\] now what do you get

OpenStudy (anonymous):

I did \[5 - 5 + 0.25 = 0.25\] what i got was simply 0.25 did i mess up??

OpenStudy (anonymous):

OK minus 5 so you want me to subratact that by 5 as well

OpenStudy (anonymous):

yes, so thats what i mean by do this to both sides

OpenStudy (anonymous):

Is that -4.75 correct? or just 0.20

OpenStudy (anonymous):

So you cannot subtract the number five from an 0.25x you just subtract it from that side Ok maybe think about it like this, \[(5-5)+0.25x = (15-5)\] what do you get

OpenStudy (anonymous):

(5- 5= 0) + 0.25x = (15 - 5 = 10) 0 + 0.25x = 10

OpenStudy (anonymous):

perfect

OpenStudy (anonymous):

so now how could you get rid of the 0.25 infront of the x you cant take it away from 10 so can you think what else you might be able to do?

OpenStudy (anonymous):

Im thinking maybe turn 10 into a decimal or add 10? or maybe i have to leave it

OpenStudy (anonymous):

So if you need to get rid of a number on its own you can do the opposite to both sides. This is what we did to get rid of the 5 so we wanted to get rid of the 5 so we took -5 because it was the opposite. Now we want to get rid of the 0.25x and end up with just x so think of it like this 0.25 times x So what is the opposite of multiplying

OpenStudy (anonymous):

division but whats x? 5?

OpenStudy (anonymous):

right division so now you have the equation:\[0.25x = 10\] right so now if we do the opposite to both sides we divide by 0.25 so this:\[\frac{ 0.25 }{ 0.25}x = \frac{ 10 }{ 0.25 }\] Do you understand?, if so what does x =

OpenStudy (anonymous):

\[I did \frac{0.25 }{10 ? } = 0.025?\]

OpenStudy (anonymous):

so its \[\frac{ 10 }{ 0.25 }\] this is not the same as\[\frac{0.25}{10}\]

OpenStudy (anonymous):

ok so the answer is 40 :o

OpenStudy (anonymous):

brilliant Also this makes sense because in your question it said that x was a number of transactions so it cant be negative and it is probably not a fraction. you can't really have half a transaction.

OpenStudy (anonymous):

So is this the answer they were looking for? no more work? thank you so much LIFE SAVER

OpenStudy (anonymous):

No problem I hope it helps